# RAG-Enhanced Collaborative LLM Agents for Drug Discovery

### v2

@@ 4.4 Ablation studies
Model Components Ablations. In Figure 2 (a), we report the results of ablations on the components of CLADD. We observe: 1) The knowledge graph and the molecular annotations are important and complementary data sources, as shown by the lower performance of the models with only Molecular Understanding or Only Knowledge Graph team available (“Only MU”, “Only KG”). 2) Dynamically selecting the relevant data sources with our Planning Team leads to better performance, leveraging their complementarity, as suggested by the lower performance of the “No Planning”. 3) The distributed architecture of the multi-agent system is a more effective way of processing the retrieved information. In “Only Planning”, we concatenate all the relevant data sources directly into the prompt of the Prediction Agent, bypassing the preprocessing and report generation of the KG and MU teams. Our results show the limitations of a single model in handling heterogeneous biological data sources. More ablations are presented in Appendix F.1.
External Knowledge Ablations. To further assess the impact of external knowledge on model performance, we evaluate the model after progressively pruning the available databases and present our results in Figure 2 (b). We observe the following: 1) Model performance depends on external knowledge size, validating the key role of the external knowledge to the framework. 2) Interestingly, we do not observe any performance plateau, indicating that further expanding the external knowledge could provide additional performance improvements. 3) From the bar plots, i.e., “No CT (No Captioning Tool)” and “Use KG (Call Knowledge Graph Team),” we observe that as the amount of external knowledge grows, the planning team increasingly depends on it. This indicates that CLADD actively leverages external knowledge more effectively during the decision-making process when such knowledge is more abundant. More analysis on external knowledge usage is given in Appendix F.2.
@@ 4.5 Case Studies
In Figure 3, we provide an example of agent collaboration in CLADD to identify the top 5 proteins a query molecule is most likely to activate. First, the BioRel Agent extracts from the knowledge graph that the anchor drug, Naftopidil, is indicated for benign prostatic hyperplasia (BPH), pointing to likely activation of pathways related to that condition. The DrugRel Agent then complements these findings by 1) connecting BPH with alpha-1 adrenergic receptors using its internal knowledge (which is confirmed in the literature (Klotsman et al., 2004)), and 2) analyzing the related drugs in the knowledge graph (e.g., Hydroxyzine and Clopamine), inferring potential interaction with histamine and dopamine receptors. Finally, the MU agent integrates these findings with its own evidence to provide a summarized report of the likely activated proteins. This example highlights the complementary nature of the agents in CLADD, where the integration of diverse information from each agent is key to its success. The distributed architecture of the model, grounded in real data sources, also leads to increased interpretability and reliability of the generated answers. Additional case studies illustrating the collaborative behavior of agents in CLADD are provided in Appendix F.3.
@@ 5 Conclusion
In this work, we introduced CLADD, a multi-agent framework for molecular question-answering that dynamically retrieves and integrates external knowledge to support various drug discovery tasks. We showcased its flexibility and effectiveness across multiple tasks, outperforming both general-purpose and domain-specific LLMs as well as standard deep learning methods, without requiring expensive domain-specific fine-tuning. Our analyses highlighted the complementarity of external knowledge sources, internal LLM reasoning, and multi-agent orchestration. CLADD’s chain of messages also provides insight into its decision-making process and the role of different data sources, fostering more interpretable scientist-AI interactions.
Future Work. Our findings revealed a strong correlation between external knowledge size and system performance, with no observed plateaus, suggesting opportunities in scaling up external data. Additionally, beyond serving as a standalone tool, CLADD could also be leveraged as a component of more complex agentic workflows, for example, combining computational and experimental systems (Tom et al., 2024), which will be the subject of future work.
@@ Appendix A Additional Related Works
LLMs with Knowledge Graphs. While large language models (LLMs) have been successfully adapted to numerous domains, they have faced criticism for their lack of factual accuracy. Specifically, LLMs often struggle to recall reliable facts and are prone to hallucinations (Ji et al., 2023), which can be a bottleneck for scientific applications. A promising approach to mitigate these issues is the integration of external knowledge sources, such as knowledge graphs (KGs), into LLMs during the generation process. For instance, Baek et al. (2023) proposes a method where relevant triplets are retrieved from KGs based on the input query. These triplets are then verbalized and provided as additional input to the LLM, enhancing its factual grounding and accuracy. KG-Rank (Yang et al., 2024) focuses on medical question-answering, leveraging a medical knowledge graph to match terms in the question and expand them. DALK (Li et al., 2024) leverages an LLM to construct an Alzheimer’s disease-specific KG, which is then used to enhance the accuracy and relevance of LLM-generated responses. Although these methods retrieve entities from KGs that are related to those in the query, the virtually infinite number of potential molecules of interest in drug discovery, combined with the limited domain expertise of general-purpose LLMs, makes it challenging to directly apply existing techniques to molecular question-answering.
@@ Appendix B Preliminaries
Tanimoto Similarity. The Tanimoto similarity is a widely accepted criterion for calculating the similarity between two molecules based on their molecular fingerprint (Bajusz et al., 2015), which are the binary sequences that denote the presence or absence of specific substructures (Rogers & Hahn, 2010). Given two molecules $g_{i}$ and $g_{j}$ with fingerprints $\mathtt{fp}_{i}$ and $\mathtt{fp}_{j}$, the Tanimoto similarity $s_{i,j}$ is computed as follows:
$\small s_{i,j}=\frac{|\mathtt{fp}_{i}\cap\mathtt{fp}_{j}|}{|\mathtt{fp}_{i}|+|\mathtt{fp}_{j}|-|\mathtt{fp}_{i}\cap\mathtt{fp}_{j}|}.$ (9)
@@ Appendix C Datasets
hERG DILI Skin Carcinogens BBBP Sider ClinTox BACE ChemPert
@@
C.1 Drug Toxicity Prediction Task
For the drug toxicity prediction task, we use four datasets: hERG, DILI, Skin, and Carcinogens.
  • The Human ether-a-go-go related gene (hERG) (Wang et al., 2016) plays a critical role in regulating the heart’s rhythm. Thus, accurately predicting hERG liability is essential in drug discovery. In this task, we assess the model’s ability to predict whether a drug blocks hERG.
  • Drug-induced liver injury (DILI) (Xu et al., 2015) is a severe liver condition caused by medications. In this task, we evaluate the model’s capability to predict whether a drug is likely to cause liver injury.
  • Repeated exposure to a chemical agent can trigger an immune response in inherently susceptible individuals, resulting in Skin (Alves et al., 2015) sensitization. In this task, we evaluate the model’s capability to predict whether the drug induces a skin reaction.
  • A Carcinogen (Lagunin et al., 2009) refers to any substance, radionuclide, or type of radiation that contributes to carcinogenesis, the development of cancer. In this task, we assess the model’s ability to predict whether a drug has carcinogenic properties.
@@ C.2 Property-Specific Molecular Captioning Task
For the property-specific molecular captioning task, we use four datasets in MoleculeNet (Wu et al., 2018): BBBP, Sider, Clintox, BACE
@@
C.3 Drug-Target Prediction Task Task
We rely on annotated molecular targets present in the Drug Repurposing Hub (Corsello et al., 2017), DrugBank (Wishart et al., 2018) and STITCH v5.0 (Szklarczyk et al., 2016), as combined and preprocessed in Zheng et al., 2023. As we explained in Section 4, we separately report the performance on the test set for molecules based on their information availability in the external databases (“Overlap”/“No Overlap”). More specifically, for “No Overlap” cases, we exclude the molecules in the following criteria:
@@ C.3 Drug-Target Prediction Task Task
After this process, 5771 molecules remained in the test set for “No Overlap” scenario.
@@ Appendix D Baselines Setup
In this section, we provide further details on the baselines we used in Section 4. For all baseline models, we utilize the pre-trained checkpoints provided by the authors of the original papers.
@@
  • Galactica (Taylor et al., 2022) is a large language model designed to store, integrate, and reason over scientific knowledge. The authors demonstrate Galactica’s capabilities in simple molecule understanding tasks, such as predicting IUPAC names and performing binary classification for molecular property prediction. We also fine-tune Galactica for the Drug-Target Prediction task described in Section 4, using molecules linked to more than five proteins for activation and inhibition. For fine-tuning, we search for the optimal hyperparameters by experimenting with learning rates of $\{1e-3,1e-4,1e-5,1e-6\}$ and epochs of $\{50,100,150,200\}$, reporting the best performance achieved.
  • GIMLET (Zhao et al., 2023a) introduces a unified approach to leveraging language models for both graph and text data. The authors aim to enhance the generalization ability of language models for molecular property prediction through instruction tuning.
  • LlaSMol (Yu et al., 2024) presents a large-scale, comprehensive, and high-quality dataset designed for instruction tuning of large language models. This dataset includes tasks such as name conversion, molecule description, property prediction, and chemical reaction prediction, and it is used to fine-tune different open-source LLMs.
@@ Appendix E Implementation Details
E.1 Software Configuration
Our model is implemented using Python 3.11, PyTorch 2.5.1, Torch-Geometric 2.6.1, RDKit 2023.9.6, and LangGraph 0.2.59.
E.2 External Databases
In all experiments, we employ the PubChem database (Kim et al., 2021) as the annotation database $\mathcal{C}$ and PrimeKG (Chandak et al., 2023) as the biological knowledge graph $\mathcal{G}$.
@@ E.2 External Databases
E.3 KG Planner
In section 3.2.1, we propose to utilize 3D geometrically pre-trained GNNs to retrieve molecules highly structurally similar to the query molecule. We use GIN architecture (Xu et al., 2018), which is pre-trained with GraphMVP (Liu et al., 2021) approach. The checkpoint of the model is available at  [^2] https://huggingface.co/chao1224/MoleculeSTM/tree/main/pretrained_GraphMVP.
@@ F.1 Additional Ablation Studies
In Table 6, we conduct a model analysis by removing one component of the model at a time. We have the following observations: 1) By comparing “Only Expert Annotation” and “Only Generated Caption”, we observe that relying solely on expert annotations yields significantly better performance. This highlights the critical importance of human-generated annotations over machine-generated captions. 2) Among the three agents—DrugRel Agent, BioRel Agent, and MU Agent—we could not determine a clear superiority in their relative importance, as the agent leading to the best performance by itself varied by task (Activation or Inhibition). 3) Overall, we observe a decline in performance when any single component of CLADD is removed, emphasizing the significance of each component.
Moreover, we perform additional ablation studies in the property-specific molecule captioning tasks in Figure 6. We observe that including all components (i.e., CLADD) leads to best performance except for the BACE dataset. This is because, as illustrated in Figure 8, the BACE dataset contains minimal relevant information in both the annotation database and the knowledge graph. Consequently, the model derives minimal benefit from external knowledge, highlighting the critical role of having relevant external information.
Table 6: Additional ablation studies in protein target tasks (Precision @ 5). Bold and underline indicate best and second-best methods.
@@
F.2 Additional External Knowledge Analysis
In Figure 7, we analyze how external knowledge is used during the decision-making process for drug-target prediction tasks. We have the following observations: 1) As shown in Figure 7 (a) and (b), the average length of human descriptions is considerably longer in the “Correct” case, and the number of retrieved 2-hop paths is notably higher in the “Correct” case. This highlights the importance of having external information that is both high quality and abundant. 2) On the other hand, although we anticipated a higher proportion of 2-hop paths containing Gene/Protein entities in the “Correct” case, no significant difference was observed between the “Correct” and “Incorrect” cases in Figure 7 (c) and (d). From these results, we argue that CLADD’s performance is not solely reliant on retrieving external information that is directly linked to the correct answer, given that external information can be further processed and contextualized by the agents, also integrating different sources of evidence.
Figure 7: External knowledge analysis results. (a) The average length of retrieved human descriptions, (b) the average number of retrieved 2-hop paths in the knowledge graph, and (c) the proportion of entity types in 2-hop paths for correct and incorrect cases.

### v3

@@ 4.4 Ablation studies
Model Components Ablations. In Figure 2 (a), we report the results of ablations on the components of CLADD. We observe: 1) The knowledge graph and the molecular annotations are important and complementary data sources, as shown by the lower performance when only Molecular Understanding or Knowledge Graph team is available (“Only MU”, “Only KG”). 2) Dynamically selecting the relevant data sources with Planning Team improves performance, leveraging their complementarity, as suggested by the lower performance of the “No Planning”. 3) The distributed architecture of the multi-agent system is a more effective way of processing the retrieved information, as highlighted by the lower performance of “Only Planning” where all the relevant data sources are directly included in the prompt of a single Prediction Agent, bypassing intermediate reports. Additional ablation studies are presented in Appendix F.1. Furthermore, we confirmed results across different LLMs, including open-source models, showcasing the LLM-agnostic nature of CLADD in Appendix F.2.
External Knowledge Ablations. To further assess the impact of external knowledge on model performance, we evaluate the model after progressively pruning the available databases and present our results in Figure 2 (b). We observe the following: 1) Model performance depends on external knowledge size, validating the key role of the external knowledge to the framework. 2) Interestingly, we do not observe any performance plateau, indicating that further expanding the external knowledge could provide additional performance improvements. 3) From the bar plots, i.e., “No CT (No Captioning Tool)” and “Use KG (Call Knowledge Graph Team)”, we observe that as the amount of external knowledge grows, the planning team increasingly depends on it. This indicates that CLADD actively leverages external knowledge more effectively during the decision-making process when such knowledge is more abundant. A more detailed analysis of how external knowledge is utilized and its impact on model performance is provided in Appendix F.3.
@@ 4.5 Case Studies
Figure 3 showcases how the agents in CLADD collaborate to identify “the top-5 protein targets a query molecule is most likely to activate”. First, the BioRel Agent extracts from the knowledge graph that the anchor drug, Naftopidil, is indicated for benign prostatic hyperplasia (BPH), implying the activation of related pathways. The DrugRel Agent complements these findings by 1) linking BPH to alpha-1 adrenergic receptors using its internal knowledge (which is confirmed in the literature (klotsman2004case)), and 2) analyzing related drugs in the knowledge graph (e.g., Hydroxyzine, Clopamine), to infer interactions with histamine and dopamine receptors. Finally, the MU agent integrates these findings with the analysis of the molecular structure to provide a summarized report of the activated protein targets. This example highlights the agents’ complementary strengths, which lead to interpretable and reliable predictions. Additional case studies are provided in Appendix F.4.
@@ 5 Conclusion
In this work, we introduced CLADD, a RAG-enhanced multi-agent framework for zero-shot molecular question-answering that can support various drug discovery tasks. We showcased its flexibility and effectiveness across multiple real-world tasks, outperforming both general-purpose and domain-specific fine-tuned LLMs. Our analyses highlighted the complementarity of external knowledge sources, internal LLM reasoning, and multi-agent orchestration. CLADD’s chain of messages also provides insight into its decision-making process, fostering more interpretable interactions.
Acknowledgement
N.L. and C.P. were supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2025-02304967, AI Star Fellowship (KAIST)). Additionally, N.L. and C.P. received funding from the National Research Foundation of Korea (NRF) through two separate grants: RS-2024-00335098 (funded by the Korea government (MSIT)) and RS-2022-NR068758 (funded by the Ministry of Science and ICT).
E.D.B., E.H., T.B., G.S. are employees of Genentech and shareholders of Roche.
@@ Appendix A Additional Related Works
LLMs with Knowledge Graphs. While large language models (LLMs) have been successfully adapted to numerous domains, they have faced criticism for their lack of factual accuracy. Specifically, LLMs often struggle to recall reliable facts and are prone to hallucinations (ji2023survey), which can be a bottleneck for scientific applications, and are still persistent after fine-tuning (gekhman-etal-2024-fine). A promising approach to mitigate these issues is the integration of external knowledge sources, such as knowledge graphs (KGs), into LLMs during the generation process. For instance, baek2023knowledge proposes a method where relevant triplets are retrieved from KGs based on the input query. These triplets are then verbalized and provided as additional input to the LLM, enhancing its factual grounding and accuracy. KG-Rank (yang2024kg) focuses on medical question-answering, leveraging a medical knowledge graph to match terms in the question and expand them. DALK (li2024dalk) leverages an LLM to construct an Alzheimer’s disease-specific KG, which is then used to enhance the accuracy and relevance of LLM-generated responses. Although these methods retrieve entities from KGs that are related to those in the query, the virtually infinite number of potential molecules of interest in drug discovery, combined with the limited domain expertise of general-purpose LLMs, makes it challenging to directly apply existing techniques to molecular question-answering.
@@ Appendix B Preliminaries
Tanimoto Similarity. The Tanimoto similarity is a widely accepted criterion for calculating the similarity between two molecules based on their molecular fingerprint (bajusz2015tanimoto), which are the binary sequences that denote the presence or absence of specific substructures (rogers2010extended). Given two molecules $g_{i}$ and $g_{j}$ with fingerprints $\mathtt{fp}_{i}$ and $\mathtt{fp}_{j}$, the Tanimoto similarity $s_{i,j}$ is computed as follows:
$s_{i,j}=\frac{|\mathtt{fp}_{i}\cap\mathtt{fp}_{j}|}{|\mathtt{fp}_{i}|+|\mathtt{fp}_{j}|-|\mathtt{fp}_{i}\cap\mathtt{fp}_{j}|}.$ (8)
@@ Appendix C Datasets
hERG DILI Skin MLSMR_Mtb BBBP Sider ClinTox BACE ChemPert
@@
C.1 Drug Biological Activity Prediction Task
For the drug biological activity prediction task, we use four datasets: hERG, DILI, Skin, and MLSMR_Mtb.
  • The Human ether-a-go-go related gene (hERG) (wang2016admet) plays a critical role in regulating the heart’s rhythm. Thus, accurately predicting hERG liability is essential in drug discovery. In this task, we assess the model’s ability to predict whether a drug blocks hERG.
  • Drug-induced liver injury (DILI) (xu2015deep) is a severe liver condition caused by medications. In this task, we evaluate the model’s capability to predict whether a drug is likely to cause liver injury.
  • Repeated exposure to a chemical agent can trigger an immune response in inherently susceptible individuals, resulting in Skin (alves2015predicting) sensitization. In this task, we evaluate the model’s ability to predict whether the drug induces a skin reaction.
  • The Molecular Libraries Small Molecule Repository - Mycobacterium tuberculosis dataset (MLSMR_Mtb) has been released as part of eke2025genetic. Antimycobacterial activity against M. tuberculosis was measured in a dose-response assay and quantified as AUC. Following the original study, we used an AUC cutoff of 25 for classification. Out of the 935 molecules tested, we randomly selected 200 compounds with a balanced positive/negative ratio. For this task, we evaluate the model’s ability to predict antimycobacterial activity. In addition to its relevance, we selected this dataset for its recency, as it was published after GPT-4o and in parallel to the preparation of this study, ensuring no overlap with pre-training data and thus allowing benchmarking against leakage risks. To the best of our knowledge, our work is the first study leveraging this dataset.
@@ C.2 Property-Specific Molecular Captioning Task
For the property-specific molecular captioning task, we use four datasets in MoleculeNet (wu2018moleculenet): BBBP, Sider, Clintox, BACE.
@@
Evaluation Protocol. While previous works on molecular captioning generate general molecule descriptions and evaluate them with standard NLP metrics like BLEU. However, because a molecule can be described in multiple ways (some more relevant to certain tasks (guo2024moltailor; edwards2024molcap)), we focus on property-specific captioning. Here, the main challenge is the lack of ground-truth captions for each property. Therefore, similar to previous work (edwards2024molcap), we use an evaluation protocol that checks how well the generated captions aid in property prediction by fine-tuning a language model (SciBERT) on them. Specifically, for a generated caption and the SMILES representation of the target molecule, we concatenate them using a [CLS] token, forming SMILES[CLS]caption, and fine-tune a SciBERT (beltagy2019scibert) model for property prediction. Importantly, fine-tuning SciBERT is only part of the evaluation protocol, as CLADD itself does not involve any fine-tuning. This process is illustrated in Figure 4.
Figure 4: (a) After CLADD (or baseline models) generates a property-specific caption, (b) SciBERT is used for evaluation. In other words, fine-tuning SciBERT is not part of CLADD; it is only used for evaluation purposes.
C.3 Drug-Target Prediction Task
We rely on annotated molecular targets present in the Drug Repurposing Hub (corsello2017drug), DrugBank (wishart2018drugbank), and STITCH v5.0 (szklarczyk2016stitch), as combined and preprocessed in zheng2023chempert. As we explained in Section 4, we separately report the performance on the test set for molecules based on their information availability in the external databases (“Overlap”/“No Overlap”). More specifically, for “No Overlap” cases, we exclude the molecules in the following criteria:
@@ C.3 Drug-Target Prediction Task
After this process, 5771 molecules remained in the test set for the “No Overlap” scenario.
@@ Appendix D Baselines Setup
This section provides further details on the baselines we used in Section 4. For all baseline models, we utilize the pre-trained checkpoints provided by the authors of the original papers.
@@
Model URL
Galactica 125M https://huggingface.co/facebook/galactica-125m
Galactica 1.3B https://huggingface.co/facebook/galactica-1.3b
Galactica 6.7B https://huggingface.co/facebook/galactica-6.7b
GIMLET https://huggingface.co/haitengzhao/gimlet
LlasMol https://huggingface.co/osunlp/LlaSMol-Mistral-7B
MolecularGPT https://huggingface.co/YuyanLiu/MolecularGPT
  • Galactica (taylor2022galactica) is a large language model designed to store, integrate, and reason over scientific knowledge. The authors demonstrate Galactica’s capabilities in simple molecule understanding tasks, such as predicting IUPAC names and performing binary classification for molecular property prediction. We also fine-tune Galactica for the Drug-Target Prediction task described in Section 4, using molecules and associated activated/inhibited proteins. For fine-tuning, we searched for the optimal hyperparameters (learning rate of $\{1e-3,1e-4,1e-5,1e-6\}$ and epoch number of $\{50,100,150,200\}$), reporting the best performance achieved.
  • GIMLET (zhao2023gimlet) introduces a unified approach to leveraging language models for both graph and text data. The authors aim to enhance the generalization ability of language models for molecular property prediction through instruction tuning.
  • LlaSMol (yu2024llasmol) presents a large-scale, comprehensive, and high-quality dataset designed for instruction tuning of large language models. This dataset includes tasks such as name conversion, molecule description, property prediction, and chemical reaction prediction, and it is used to fine-tune different open-source LLMs.
@@ Appendix E Implementation Details
Software Configuration. Our model is implemented using Python 3.11, PyTorch 2.5.1, Torch-Geometric 2.6.1, RDKit 2023.9.6, and LangGraph 0.2.59.
Computational Resources. For LLMs, we utilize the OpenAI API, thereby leveraging OpenAI’s computational resources. All other computations, such as GNN retrievers, are performed on a 24GB NVIDIA GeForce RTX 3090 GPU.
External Databases. In all experiments, we employ the PubChem database (kim2021pubchem) as the annotation database $\mathcal{C}$ and PrimeKG (chandak2023building) as the biological knowledge graph $\mathcal{G}$.
@@
E.1 KG Planner
As explained in section 3.1, we utilize a pre-trained GNN (with 3D information) to retrieve molecules highly related to the query molecule. In particular, the model has a GIN architecture (xu2018powerful), which is pre-trained with the GraphMVP (liu2021pre) approach. The checkpoint of the model is available at https://huggingface.co/chao1224/MoleculeSTM/tree/main/pretrained˙GraphMVP.
@@ F.1 Additional Ablation Studies
In Table 7, we conduct a model analysis by removing one component of the model at a time for the drug-target prediction task. We have the following observations: 1) By comparing “Only Expert Annotation” and “Only Generated Caption”, we observe that relying solely on expert annotations yields significantly better performance. This highlights the critical importance of human-generated annotations over machine-generated captions. Still, their combination leads to the best overall performance. 2) Among the three agents—DrugRel Agent, BioRel Agent, and MU Agent—we could not determine a clear superiority in their relative importance, as it was task-dependent (Activation or Inhibition). 3) Overall, we observe a decline in performance when any single component of CLADD is removed, emphasizing the significance of each module.
We perform additional ablation studies in the property-specific molecule captioning task in Figure 7. Similarly, we observe that including all components (i.e., CLADD) leads to the best performance except for the BACE dataset. Our analysis showed that this is because, as illustrated in Figure 11, the BACE dataset contains minimal relevant information in both the annotation database and the knowledge graph. Consequently, the model derives minimal benefit from external knowledge, highlighting the critical role of having relevant external information to boost performance.
Table 6: Additional ablation studies in drug-target prediction task (Precision @ 5). Bold and underline indicate best and second-best methods.
@@
F.2 LLM-Agnostic Nature of CLADD
Due to the expensive API costs, we mainly report the results using GPT-4o mini in the main manuscript to validate the proposed framework. In this section, we performed additional experiments replacing it with different LLMs, including Llama3.3-70b and DeepSeek-V3. As shown in Table 7, the proposed framework (+ CLADD) consistently improves each individual LLM, showcasing its LLM-agnostic advantage.
Table 7: Performance of CLADD with each agent replaced by a different closed-source and open-source LLM (drug-target prediction task).
Overlap No overlap
Activate Inhibit Activate Inhibit
GPT-4o mini 1.15 1.02 1.13 0.87
+ CLADD 3.04 4.83 2.67 3.24
Llama-3.3-70B 0.84 0.94 0.88 0.81
+ CLADD 3.13 6.40 2.73 4.14
DeepSeek-V3 1.91 1.46 1.90 1.11
+ CLADD 3.60 7.75 3.15 5.01
F.3 Additional External Knowledge Analysis
In Table 9, we analyze how the retrieval accuracy affects the model performance. To do so, we investigated two settings: one where the anchor drug selection in the knowledge graph is done randomly, and another where annotations are randomly sampled from the annotation database. As expected, we observe that the performance of both these models is significantly lower compared to the original model. We also observe that there is still a significant performance gap when compared to GPT-4o mini. This is expected, as our model still includes a planning team that ensures that the anchor drug and annotations are only used when they are relevant to the query molecule and task.
Moreover, we further investigate how the quality of retrieved knowledge affects the model performance. Firstly, we analyzed how performance changes as a function of the length of the annotation retrieved from the annotation database. In Figure 9(a), “Zero” indicates that no annotation is available in the annotation database, while Q1, Q2, Q3, and Q4 represent the quartiles of the retrieved annotation length. We highlight two interesting trends: (1) in general, performance increases with the annotation length, which is in line with the intuition that longer annotations include more relevant information, and (2) on average, “no annotation” leads to better results than the shortest annotations, which could indicate that the shortest annotations are often not informative enough to boost performance. However, for all groups except the shortest annotations, the additional information provides a proportional improvement.
[…]
Figure 8: Performance analysis on retrieval errors.
Overlap No overlap
Activate Inhibit Activate Inhibit
GPT-4o mini 1.15 1.02 1.13 0.87
Random Anchor Drug in KG 2.49 4.46 1.86 2.31
Random Annotations in DB 2.62 4.08 2.51 2.85
CLADD 3.04 4.83 2.67 3.24
Figure 9: Performance as a function of (a) the length of the text retrieved from the annotation database and (b) the Tanimoto similarity between the anchor molecule and the knowledge graph.
In Figure 10, we analyze how external knowledge is used during the decision-making process for the drug-target prediction task. We have the following observations: 1) As shown in Figures 10(a) and 10(b), the average length of human descriptions is considerably longer in the “Correct” case, and the number of retrieved 2-hop paths is notably higher in the “Correct” case. This highlights the importance of having external information that is both high quality and abundant. 2) On the other hand, although we anticipated a higher proportion of 2-hop paths containing Gene/Protein entities in the “Correct” case, no significant difference was observed between the “Correct” and “Incorrect” cases in Figures 10(c) and 10(d). From these results, we argue that CLADD’s performance is not solely reliant on retrieving external information that is directly linked to the correct answer, given that external information can be further processed and contextualized by the agents, integrating different sources of evidence and internal knowledge.
Figure 10: External knowledge analysis results. (a) The average length of retrieved human descriptions, (b) the average number of retrieved 2-hop paths in the knowledge graph, and (c-d) the proportion of entity types in 2-hop paths for correct and incorrect cases.